OpenStudy (anonymous):

Use the substitution method for 2x-y=3 7x-3y=11 what's the ordered pair and the solution?

OpenStudy (anonymous):

for example solve first one for y and insert the whole thing into the second y

OpenStudy (anonymous):

I have tried that but my answers are not coming out to be 3 and 11 when I do the check

jhonyy9 (jhonyy9):

@phanta_seea so you have thought it right solve it please for medal

OpenStudy (anonymous):

2x-y=3 2x=3+y ---------- y=2x-3 7x-3(2x-3)=11 7x-6x+9=11 x=2

OpenStudy (anonymous):

I got 2 for the first equation and 3 for the next one. All these steps really confuse me am I on the right

OpenStudy (anonymous):

y=2x-3 if you put x=2 you get y=4-3 y=1

OpenStudy (anonymous):

okay I see what I did wrong, thanks!!

OpenStudy (anonymous):

cool, no problem :)